# Electromagnetic $|G_{E}/G_{M}|$ ratios of hyperons at large timelike $q^{2}$

### v2

@@ Abstract
In recent years, it has become possible to measure not only the magnitude of the electric ($G_{E}$) and magnetic ($G_{M}$) form factors of spin $\frac{1}{2}$ baryons, but also to measure the relative phases of those quantities in the timelike kinematic region. Aiming to interpret present $|G_{E}/G_{M}|$ data on hyperons of the baryon octet, as well as to predict future data, we present model calculations of that ratio for large invariant 4-momentum square $q^{2}$ in the timelike region ($q^{2}>0$). Without any further parameter fitting, we extend to the timelike region a covariant quark model previously developed to describe the kinematic spacelike region ($q^{2}\leq 0$) of the baryon octet form factors. The model takes into account both the effects of valence quarks and the excitations of the meson cloud which dresses the baryons. We report here that our results are in good agreement with very recent data from BESIII, within the uncertainty limits of our model. We also make new predictions which may be tested in the near future, when the range of the present measurements is expanded to the 20–50 GeV2 region.
preprint: LFTC-24-08/91
I Introduction
@@
Although the initial measurements were performed near threshold, it is expected that with increased accelerator power, the experimental energy will go higher up to the 20–50 GeV2 range. This extension will allow us to test the predictions described in this work.
Measurements of the effective elastic form factors of the $\Lambda$, $\Sigma^{0,\pm}$, $\Xi^{0,-}$ as well as the transition form factors $\Sigma^{0}\to\Lambda$ have been performed at BaBar [17], CLEO [18, 3], Belle [19] and BESIII [20, 21, 16, 22, 4, 31, 5, 24, 30, 29, 23, 28, 5, 25, 26, 27, 24]. Measurements of the ratios $|G_{E}/G_{M}|$ for the $\Lambda$ and $\Sigma^{+}$ have also been determined recently at BESIII [16, 4, 5]. Complete experiments, that measure $|G_{E}|$, $|G_{M}|$, and the relative phase between $G_{E}$ and $G_{M}$ have been performed for the $\Lambda$ and $\Sigma^{+}$ at BESIII [4, 5], and even more measurements are planned for a near future at BESIII and PANDA [32, 33].
@@
The numerical calculations are based on the covariant spectator quark model formalism for the spacelike region [7, 34, 36, 35] extended to the timelike region. The model has been used in the study of electromagnetic structure of nucleon resonances [36, 37, 38, 39, 40, 41] and other baryons systems [43, 44, 42].
@@ II Formalism
Finally, at threshold, where $q^{2}=4M_{B}^{2}$, the functions $G_{E}$ and $G_{M}$ are such that $G_{E}=G_{M}$, as a consequence of the definition of these Sachs from factors in terms of the kinematic independent Dirac and Pauli form factors. Therefore one expects that $G_{E}/G_{M}\simeq 1$ for measurements near threshold, and the non-zero phases between the form factors are to be sought above the threshold only ($q^{2}>4M_{B}^{2}$[2].
@@
III Calculation of electromagnetic form factors at large $q^{2}$
@@
Due to the character of this approximation, we expect our estimates here to be accurate only for large $q^{2}$, deviating from the condition $G_{E}=G_{M}$ at the threshold.
Since the relations (4) are valid for the large-$|q^{2}|$ mathematical limit, and we want to make predictions to finite $q^{2}$, we consider finite corrections to these asymptotic relations. We start noticing that the elastic form factors are divided into two regions of $q^{2}$: $(-\infty,0]$ (spacelike region) and $[4M_{B}^{2},+\infty)$, with a gap of $4M_{B}^{2}$ length between the photon point ($q^{2}=0$) and the threshold of the baryon-antibaryon production ($q^{2}=4M_{B}^{2}$). The interval $(0,4M_{B}^{2})$ defines an unphysical region where no physical baryons can be detected [10, 11].
There is then an ambiguity about the center of reflection for the asymptotic symmetry relation of Eq. (4). To take into account the ambiguity of the location of the center of symmetry we replace $q^{2}\to q^{2}\left(1-\frac{2M_{B}^{2}}{q^{2}}\right)$ in (4), which introduces finite corrections ($-2M_{B}^{2}$) to the limit $q^{2}=+\infty$. Therefore we use [6]
@@
[…]
@@
For a control of the approximation we explored the band of variation of the form factors with respect to a reasonable variation of the symmetry point around this central value: when the functions $G_{\ell}(q^{2})$ are positive (negative), since they are expected to monotonically decrease (increase) with increasing $q^{2}$, we take the upper (lower) limit values to be $G_{\ell}(q^{2})=G_{\ell}^{\rm SL}(q^{2}-4M_{B}^{2})$ and the lower (upper) limit $G_{\ell}(q^{2})=G_{\ell}^{\rm SL}(q^{2})$. This procedure provides a band of variation of the finite corrections to the asymptotic symmetry identity. When $q^{2}$ is very large, the variation of the results becomes very narrow and the three estimates (central, upper and lower limits) converge to the same value, providing accurate predictions that can be tested by experimental data. Moreover, the comparison with data in that region provides information of the onset of the regime of the asymptotic symmetry (4).
IV Numerical results for $|G_{E}/G_{M}|$
@@
Our calculations of the ratio $|G_{E}/G_{M}|$ and their comparison with the data are shown in Figs. 3 and 3, that depict the results for $\Lambda$, $\Xi^{-,0}$ and the $\Sigma^{0,\pm}$, respectively. In the graphs for the $\Lambda$ and $\Sigma^{+}$ we include the available data for $|G_{E}/G_{M}|$. How we estimate the uncertainty of the results of our model for each form factor was defined in the previous section. The bands of variation for the magnitude of $G_{E}/G_{M}$ are determined from the independent band intervals of variation for the $G_{E}$ and $G_{M}$ functions.
The results and their uncertainty bands near the threshold provide only qualitative estimates since our results are expected to be valid only for large $q^{2}$, as discussed in the previous section. The deviation from the result $|G_{E}/G_{M}|=1$ at threshold is a consequence of the approximation used. The signs of our results for the form factors in the asymptotic large $q^{2}$ values are indicated on the right side in Figs. 3 and 3.
@@
which is a fraction of the measured value to $|G_{E}/G_{M}|$. This does not only show that model calculations of $|G_{E}/G_{M}|$ based on real form factors agree with the absolute value of the real part of $G_{E}/G_{M}$ when the imaginary parts are negligible. The result (7) allows also us to infer from the comparison between the experimental data for the r.h.s. and our model results for the ratio $\frac{|G_{E}|}{|G_{M}|}$ (where $G_{E}$ and $G_{M}$ are real functions) if the regime where the imaginary parts of the form factors vanish has been attained. The comparison of our model results for $|G_{E}/G_{M}|$ with the experimental values for $|{\rm Re}(G_{E}/G_{M})|$ from Refs. [16, 31] that measure phases, are displayed in Fig. 3. In the graphs we notice the agreement of our results for the form factors (considering the model uncertainty band) and the experimental data point at the largest $q^{2}$ point measured, $q^{2}\simeq 14.3$ GeV2 for the $\Lambda$ and $q^{2}\simeq 8.4$ GeV2 for the $\Sigma^{+}$. In the graphs we ignore the uncertainties of the angles $\Delta\Phi$, for simplicity, and used their central value. The consideration of the uncertainties on $\Delta\Phi$ only increases the error bars of the $|{\rm Re}(G_{E}/G_{M})|$ data. Hopefully, future experiments will be more precise, which will help to determine the point were our model calculations provide a good description of the data and determine the large $q^{2}$ region where reflection symmetry sets in.
V Effective form factors
In a previous work [6] we presented predictions for the effective form factors (1) for the baryon octet members $\Lambda$, $\Sigma^{0,\pm}$ and $\Xi^{0,-}$ for $q^{2}=10$–60 GeV2. At the time, only $\Lambda$, $\Sigma^{+}$, $\Sigma^{-}$, $\Xi^{0}$ and $\Xi^{-}$ could be tested in the region of 12–15 GeV2 by the data from CLEO [3]. Our estimates were on average consistent with the high $q^{2}$ data. Since then more data for $|G(q^{2})|$ became available. In particular data for $\Sigma^{+}$ and $\Sigma^{-}$ became available from BESIII [28] for the range 5.7–9.1 GeV2. Surprisingly, the data are in excellent agreement with our estimates (namely with the central line values, see Supplementary Material [85]) even outside the region where model validity was expected to be justified. A rough estimate of the lower limit for the application of our model based on the asymptotic relations, can be $q^{2}\approx 8M_{B}^{2}$ or $q^{2}\approx 10$ GeV2.
@@ V Effective form factors
It is clear from the figures that there is an excellent agreement between our predictions and the data, indicating that the range of the asymptotic form for the effective form factors may not be too far from the present day experimentally studied region. Future data for other hyperons may confirm or deny this tendency.
Refer to caption           Refer to caption
Refer to caption           Refer to caption
@@ VI Outlook and conclusions
In the last few years it became possible to determine $|G_{E}|$ and $|G_{M}|$ and the relative phase between them for $\Lambda$ and $\Sigma^{+}$. We expect that the ratio $|G_{E}/G_{M}|$ became accessible in the following years for most ground state hyperons ($\Sigma^{0,-}$ and $\Xi^{0,-}$), as well as for charmed baryons (starting with $\Lambda_{c}^{+}$). Our calculations show a partial agreement with the data in the range $q^{2}\simeq 14$ GeV2 ($\sqrt{s}\simeq 3.8$ GeV), which may be an indication that our model calculations may be compared with future measurements in the range of $q^{2}=20$–30 GeV2 ($\sqrt{s}=4.5$–5.5 GeV) or higher. Also the separation between the real and imaginary components may be tested in that range of energies.
Acknowledgements.

### v3

@@ Abstract
In recent years, it has become possible to measure not only the magnitude of the electric ($G_{E}$) and magnetic ($G_{M}$) form factors of spin $\frac{1}{2}$ baryons, but also to measure the relative phases of those quantities in the timelike kinematic region. Aiming to interpret present $|G_{E}/G_{M}|$ data on hyperons of the baryon octet, as well as to predict future data, we present model calculations of that ratio for large invariant 4-momentum square $q^{2}$ in the timelike region ($q^{2}>0$). Without any further parameter fitting,we extend to the timelike region a covariant quark model previously developed to describe the kinematic spacelike region ($q^{2}\leq 0$) of the baryon octet form factors. The model takes into account both the effects of valence quarks and the excitations of the meson cloud which dresses the baryons. This application to the timelike region assumes an approximation based on unitarity and analyticity that is valid only in the large $q^{2}$ region. Using the recent data from BESIII we establish the regime of validity of this approximation. We report here that our results for the effective form factor (combination of $|G_{E}|$ and $|G_{M}|$) are in good agreement with the data already for $q^{2}$ values above 15 GeV2. In addition, a more conservative onset of the validity of the approximation is provided by the newly available $|G_{E}/G_{M}|$ data which suggest that our predictions may be compared against data for $q^{2}\geq$ 20 GeV2. This is expected in the near future, when the range of the present measurements is expanded to the 20–50 GeV2 region.
journal: Physics Letters B
1 Introduction
@@
Although the initial measurements were performed near the threshold, it is expected that with increased accelerator power, the experimental energy will go higher up to the 20–50 GeV2 range. This extension will allow us to test the predictions described in this work.
Measurements of the effective elastic form factors of the $\Lambda$, $\Sigma^{0,\pm}$, $\Xi^{0,-}$ as well as the transition form factors $\Sigma^{0}\to\Lambda$ have been performed at BaBar [17], CLEO [3, 18], Belle [19] and BESIII [4, 5, 16, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Measurements of the ratios $|G_{E}/G_{M}|$ for the $\Lambda$ and $\Sigma^{+}$ have also been determined recently at BESIII [4, 5, 16]. Complete experiments, that measure $|G_{E}|$, $|G_{M}|$, and the relative phase between $G_{E}$ and $G_{M}$ have been performed for the $\Lambda$ and $\Sigma^{+}$ at BESIII [4, 5], and even more measurements are planned for a near future at BESIII and PANDA [32, 33].
@@
The numerical calculations are based on the covariant spectator quark model formalism for the spacelike region [7, 34, 35, 36] extended to the timelike region using asymptotic relations valid for large $q^{2}$ [6]. The model has been used in the study of electromagnetic structure of nucleon resonances [36, 37, 38, 39, 40, 41] and other baryons systems [42, 43, 44].
@@ 2 Formalism
Finally, at threshold, where $q^{2}=4M_{B}^{2}$, the functions $G_{E}$ and $G_{M}$ are such that $G_{E}=G_{M}$, as a consequence of the definition of these Sachs from factors in terms of the kinematic independent Dirac and Pauli form factors. Therefore one expects that $G_{E}/G_{M}\simeq 1$ for measurements near the threshold, and the non-zero phases between the form factors are to be sought above the threshold only ($q^{2}>4M_{B}^{2}$[2].
@@
3 Calculation of electromagnetic form factors at large $q^{2}$
@@
The relations (4) are strictly valid in the mathematical limit $q^{2}\to+\infty$. Alone the relations provide no information on the range where approximated results can be expected to have a certain accuracy. But the large physical $q^{2}$ scale where the mathematical theorem starts to be valid can be determined by the comparison with the physical data, which we do here.
@@
Given the character of approximation, based on asymptotic relations, we expect our estimates here to be accurate only for large $q^{2}$, deviating from the condition $G_{E}=G_{M}$ at the threshold.
Since the relations (4) are valid for the large-$|q^{2}|$, and we want to make predictions for finite $q^{2}$, we consider corrections to these asymptotic relations. We start by noticing that the elastic form factors are divided into two regions of $q^{2}$: $(-\infty,0]$ (spacelike region) and $[4M_{B}^{2},+\infty)$, with a gap of $4M_{B}^{2}$ length between the photon point ($q^{2}=0$) and the threshold of the baryon-antibaryon production ($q^{2}=4M_{B}^{2}$). The interval $(0,4M_{B}^{2})$ defines an unphysical region where no physical baryons can be detected [10, 11], but where the peak structure in the form factors can be inferred from the $e^{+}e^{-}\to\pi^{+}\pi^{-}$ cross sections. Only for much larger values of $q^{2}$ above the threshold ($q^{2}=4M_{B}^{2}$), we expect that the tail of the form factors becomes symmetric to the smooth one of the spacelike region ($q^{2}\leq 0$).
It is then clear that $q^{2}=0$ is not the center of reflection of the asymptotic symmetry relation (4). But there is an ambiguity about the exact location of this center, and to take into account this ambiguity we replace $q^{2}\to q^{2}\left(1-\frac{2M_{B}^{2}}{q^{2}}\right)$ in (4), which introduces finite corrections ($-2M_{B}^{2}$) to the limit $q^{2}=+\infty$. Therefore we use [6]
@@
for $\ell=E,M$. The label SL on the r.h.s. indicates that the $G_{\ell}^{\rm SL}(Q^{2})$ is calculated in the spacelike region ($Q^{2}=-q^{2}\geq 0$). The expression (5) estimates the reflection symmetry point underlying (4) exactly at the center of the unphysical interval $(0,4M_{B}^{2})$.
@@
For a control of the approximation we explored the band of variation of the form factors with respect to a reasonable variation of the symmetry point around this central value: when the functions $G_{\ell}(q^{2})$ are positive (negative), since they are expected to monotonically decrease (increase) with increasing $q^{2}$, we take the upper (lower) limit values to be $G_{\ell}(q^{2})=G_{\ell}^{\rm SL}(q^{2}-4M_{B}^{2})$ and the lower (upper) limit $G_{\ell}(q^{2})=G_{\ell}^{\rm SL}(q^{2})$. This procedure provides a band of variation of the finite corrections to the asymptotic symmetry identity. When $q^{2}$ is very large, the variation of the results becomes very narrow and the three estimates (central, upper and lower limits) converge to the same value, providing accurate predictions that can be tested by experimental data.
The relations (5) are naturally valid in the perturbative QCD (pQCD) region, where the form factors are regulated by power laws [81, 82]. Is it possible, however, that they start to be observed earlier than the pQCD regime. The origin of the symmetry relations is unitarity and analyticity. To begin with, the hadron bound states and decays of the hadrons on meson-baryon states seen as peaked structures of the timelike form factors are limited to the confinement region (lower energy). The comparison with data provides information on the onset of the regime of the asymptotic reflection symmetry (4). This is further discussed in Section 5 on the effective form factors.
4 Numerical results for $|G_{E}/G_{M}|$
@@
Our calculations of the ratio $|G_{E}/G_{M}|$ and their comparison with the data are shown in Figs. 1 and 2, that depict the results for $\Lambda$, $\Xi^{-,0}$ and the $\Sigma^{0,\pm}$, respectively. In the graphs for the $\Lambda$ and $\Sigma^{+}$ we include the available data for $|G_{E}/G_{M}|$. How we estimate the uncertainty of the results of our model for each form factor was defined in the previous section. The bands of variation for the magnitude of $G_{E}/G_{M}$ are determined from the independent band intervals of variation for the $G_{E}$ and $G_{M}$ functions.
The results and their uncertainty bands near the threshold provide only qualitative estimates since our results are expected to be valid only for large $q^{2}$, as discussed in the previous section. The deviation from the result $|G_{E}/G_{M}|=1$ at threshold is a consequence of the approximation used. The signs of our results for the form factors in the asymptotic large $q^{2}$ values are indicated on the right side in Figs. 1 and 2.
@@
Refer to caption           Refer to caption
Refer to caption           Refer to caption
@@
which is a fraction of the measured value to $|G_{E}/G_{M}|$. This does not only show that model calculations of $|G_{E}/G_{M}|$ based on real form factors agree with the absolute value of the real part of $G_{E}/G_{M}$ when the imaginary parts are negligible. The result (7) allows also us to infer from the comparison between the experimental data for the r.h.s. and our model results for the ratio $\frac{|G_{E}|}{|G_{M}|}$ (where $G_{E}$ and $G_{M}$ are real functions) if the regime where the imaginary parts of the form factors vanish has been attained. The comparison of our model results for $|G_{E}/G_{M}|$ with the experimental values for $|{\rm Re}(G_{E}/G_{M})|$ from Refs. [16, 31] that measure phases, are displayed in Fig. 3. In the graphs we notice the agreement of our results for the form factors (considering the model uncertainty band) and the experimental data point at the largest $q^{2}$ point measured, $q^{2}\simeq 14.3$ GeV2 for the $\Lambda$ and $q^{2}\simeq 8.4$ GeV2 for the $\Sigma^{+}$. In the graphs we ignore the uncertainties of the angles $\Delta\Phi$, for simplicity, and used their central value. The consideration of the uncertainties on $\Delta\Phi$ only increases the error bars of the $|{\rm Re}(G_{E}/G_{M})|$ data.
It is worth noticing that the agreement between model calculations and the data is not an indication that we are in the pQCD regime. Evidences of scaling with the pQCD behavior for the nucleon are observed only for $Q^{2}\geq 30$ GeV2 [82]. One expects, however, that the convergence for the pQCD regime can be seen for lower values of $q^{2}$ for hyperons due to dynamics of the strange quarks.
Future experiments with more precise data can help to determine the point where our model calculations provide a good description of the data and determine the $q^{2}$ region where reflection symmetry sets in.
5 Effective form factors
In a previous work [6] we presented predictions for the effective form factors (1) for the baryon octet members $\Lambda$, $\Sigma^{0,\pm}$ and $\Xi^{0,-}$ for $q^{2}=10$–60 GeV2. At the time, only $\Lambda$, $\Sigma^{+}$, $\Sigma^{-}$, $\Xi^{0}$ and $\Xi^{-}$ could be tested in the region of 12–15 GeV2 by the data from CLEO [3]. Our estimates were on average consistent with the high $q^{2}$ data. Since then more data for $|G(q^{2})|$ became available. In particular data for $\Sigma^{+}$ and $\Sigma^{-}$ became available from BESIII [28] for the range 5.7–9.1 GeV2. Surprisingly, the data are in excellent agreement with our estimates (namely with the central line values, see Supplementary Material [87]) even outside the region where model validity was expected to be justified. A rough estimate of the lower limit for the application of our model based on the asymptotic relations, can be $q^{2}\approx 10$ GeV2 or $q^{2}\approx 8M_{B}^{2}$.
@@ 5 Effective form factors
It is clear from the figures that there is an excellent agreement between our predictions and the data for $q^{2}>15$ GeV2, indicating that we are already in the range of the asymptotic form for the effective form factors. Compared to the convergence to the pQCD regime, a faster convergence of the reflection symmetry relations to their asymptotic behavior can be expected because in the effective form factors $G_{M}$ dominates on $|G|$ for large $q^{2}$: for instance, for the case of the nucleon, $G_{M}$ dominates over $G_{E}$, and $G_{M}$ converges to the expected power law faster than $G_{E}$, in the spacelike region [34, 44].
Future data may confirm or deny if we are also close to the range of the asymptotic region for the ratio $|G_{E}/G_{M}|$.
@@ 6 Outlook and conclusions
In the last few years it became possible to determine $|G_{E}|$ and $|G_{M}|$ and the relative phase between them for $\Lambda$ and $\Sigma^{+}$. We expect that the ratio $|G_{E}/G_{M}|$ became accessible in the following years for most ground state hyperons ($\Sigma^{0,-}$ and $\Xi^{0,-}$), as well as for charmed baryons (starting with $\Lambda_{c}^{+}$). Our calculations show the start of a partial agreement with the data for $q^{2}\simeq 14$ GeV2 ($\sqrt{s}\simeq 3.8$ GeV), which may be an indication that our model calculations may be compared with future measurements in the range of $q^{2}=20$–30 GeV2 ($\sqrt{s}=4.5$–5.5 GeV) or higher. Also the separation between the real and imaginary components may be tested in that range of energies.
Acknowledgments